# Accelerating Equity: Overcoming the Gender Gap in VC Funding

### v1

@@ 5 Results
Since the first stage is not the primary focus of our analysis, we describe its estimation routine in Appendix A.5, while Appendix A.6 reports the estimation result. The model fits the data well, explaining 78% of the observed and unobserved variation in match qualities. In this section, instead, we focus on the second-stage estimates and the gender gap.
@@ 5.1 Post-Accelerator Performance
The first four columns of Table 6 present the OLS regression results for startup VC fundraising performance one year after the demo day. [^15] Across tables we always include the set of regressors discussed in Section 4.3. The independent variables in the regression include founder characteristics such as Female Founder, Female Founder $\times$ ”All Men” Accelerator, No Serial Founder, At least One Graduate Degree, At least One PhD Degree, At least One Engineering/Science Degree, Average Age of Founding Team, and Founding Team Size. Additionally, startup and accelerator characteristics include Startup Age, Startup Relocated, Accelerator in Startup Hubs (CA, NY, MA), Accelerator Experiences (years), Log(Cohort Size), and ”All Men” Accelerator. The regressions also control for, the Correction Term ($\mathbb{E}[\eta^{as}|\mu,X]$), Year Fixed Effects, and Industry Fixed Effects. We begin with the one-year horizon because such funding is typically initiated during or shortly after the demo day, as the process from signing the initial contract to receiving the investment often takes several months. [^16] https://www.forbes.com/sites/alejandrocremades/2019/01/03/how-long-it-takes-to-raise-capital-for-a-startup/?sh=1129c4a37a41 Within this short period, the quality of startups changes minimally following graduation, providing us with sufficient control over the startup qualities that are observable to accelerators and investors but not to researchers.
@@
All models control for graduation year effects and include a correction term for unobserved startup-accelerator match quality from Proposition 1. The results in columns (1) and (2) show that the coefficients on “Female Founder” are not statistically significant, suggesting no significant gender differences in the likelihood of receiving any VC funding or amounts below two million US\$. However, in columns (3) and (4), the coefficients on “Female Founder” become significantly negative, indicating that women-founded startups are significantly less likely to raise larger amounts of funding.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Funding Within One Year Since Graduation Funding Within Five Years Since Graduation Funded? Funding Exceeds (0/1) Funded? Funding Exceeds (0/1) (0/1) One Million Two Millions Five Millions (0/1) One Million Two Millions Five Millions Female Founder 0.053 -0.049 -0.082*** -0.038* -0.029 -0.156 -0.109 -0.160** (0.130) (0.102) (0.029) (0.021) (0.123) (0.124) (0.116) (0.073) "All Men" Accelerator 0.049 0.051 -0.023 -0.009 0.085 0.055 0.051 0.007 (0.062) (0.052) (0.034) (0.025) (0.064) (0.067) (0.061) (0.053) Female Founder $\times$ "All Men" Accelerator -0.097 0.068 0.066* 0.017 -0.097 0.168 0.068 0.103 (0.149) (0.113) (0.039) (0.024) (0.145) (0.143) (0.133) (0.090) Startup Relocated -0.068 -0.105*** -0.043** -0.033** -0.046 -0.084 -0.081* -0.034 (0.052) (0.038) (0.022) (0.016) (0.057) (0.053) (0.045) (0.042) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $N$ 736 736 736 736 736 736 736 736 $R^{2}$ 0.093 0.117 0.055 0.034 0.097 0.087 0.056 0.048 * – $p<0.1$; ** – $p<0.05$; *** – $p<0.01$
  • Notes: This table presents linear probability estimates from (3), where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
@@
Across columns, we also examine whether “all-man” accelerators further decrease the chances of female entrepreneurs obtaining funding, as prior studies have shown that female investors do not display bias against female entrepreneurs and can help mitigate the gender gap (Raina,, 2021; Hebert,, 2023). The coefficients of “Female Founder $\times$×All Men’ Accelerator” are generally not significant for the probability of being funded (columns 1 and 5). However, it is surprisingly positive and significantive in column (3) for the probability of receiving funding of more than 2 million US\$. Taken together, this evidence does not support the idea that startups with women founders are more likely to match with accelerators with a mixed (or all-women) gender composition.
Finally, the last row of Table 9 highlights the significant costs associated with relocation. Moving to a new state reduces the probability of reaching the one-million and five-million funding milestones by approximately 3.3 to 10 percentage points. To the extent that startups securing five million in funding are of higher quality, their superior capabilities may help them better absorb these relocation costs. Over the five-year horizon, however, the negative impact of relocation diminishes substantially.
@@ 6 What Drives the Gender Gap in Funding?
In Table 7, we include the interaction of “Female Founder” and “Startup Relocated.” As in the previous table, the first four columns report startups’ performance one year after graduation, while the remaining columns extend the analysis to five years. Importantly, because of our control function approach (Proposition 1) we can consider startup relocation as exogenous as relocation was controlled for in the matching utilities (1)—as for all the other covariates in (3)—ensuring a causal interpretation of our findings.
Across columns, the estimated coefficients for “Female Founder” closely align with those in Table 6. For example, female-founded teams have a 7.6 percentage point lower probability of securing at least two million US\$ in the first year after graduation and a 17.9 percentage point lower probability of raising five million US\$ within five years. However, this gap narrows for startups willing to relocate.
The second row presents the estimates for the interaction term. Despite large standard errors due to the small sample, the coefficients are generally positive, suggesting that relocation mitigates the gender gap in funding.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Funding Within One Year Funding Within Five Years Funded? Funding Exceeds (0/1) Funded? Funding Exceeds (0/1) (0/1) One Mil. Two Mil. Five Mil. (0/1) One Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.014 -0.078 -0.076** -0.041* -0.087 -0.195 -0.111 -0.179** (0.135) (0.101) (0.030) (0.023) (0.132) (0.129) (0.126) (0.086) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.238* 0.102 -0.022 0.010 0.209 0.139 0.007 0.065 (0.143) (0.116) (0.034) (0.019) (0.143) (0.146) (0.129) (0.106) Other Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$ 0.225 0.024 -0.098** -0.031 0.122 -0.056 -0.104 -0.113 (0.149) (0.150) (0.039) (0.029) (0.144) (0.164) (0.140) (0.103) $N$ 736 736 736 736 736 736 736 736 $R^{2}$ 0.097 0.118 0.055 0.034 0.100 0.088 0.056 0.048 * – $p<0.1$; ** – $p<0.05$; *** – $p<0.01$
  • Notes: This table presents linear probability estimates from (3) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
Higher-quality startups may be more likely to relocate, raising the concern that our analysis captures unobserved aspects of startup quality correlated with female-founded teams rather than the actual match quality and complementarity between startups and accelerators.
@@ 6.1 Mechanism
Costs for mother founders. To understand the mechanism, we replicate the analysis above, focusing on the probability of observing fundings exceeding two and five million US\$ at either horizon, as the previous analyses highlight that the gender gap is especially pronounced for these funding levels.
Relocation may be particularly costly for mother entrepreneurs. Relocating with a newborn or a young child may be particularly difficult, regardless of the quality of the startup. To understand whether the gender gap is particularly severe in these cases, we subset the data to include only startups whose teams’ average age is between 28 and 40 years old. We chose this range because there is evidence that women with a degree tend to have their first child later than other mothers, with a median age of first childbirth at 28. This age further increases for women in STEM fields, as is relevant to our sample of startup founders (Livingston,, 2015; Lappegard et al.,, 2020; Schweizer and Guzzo,, 2020).
The results in Table 8 indicate a larger gender gap (row 1) compared to the estimates obtained from the full sample. The gender gap appears particularly large over the five-year horizon (columns 3 and 4). For instance, female founders who do not relocate are 34 percentage points less likely to reach both the two- (non-significant) and the five-million funding milestones. This result comes as no surprise, as parenting may take significant resources from mother entrepreneurs, and the time they allocate away from their startups has a substantial economic cost five years after graduation.
However, the second line provides new insights into the benefits of relocation. While there seems to be no advantage over short horizons (columns 1 and 2), there are substantial gains from (successfully) relocating over the five-year horizon. Here, relocating female founders improve their odds of reaching the two-million milestone by 26.5 percentage points and the five-million milestone by 41 percentage points. We show that this effect compensates for the gender gap, opening up possibilities for policy solutions that support young women in relocating. We turn to what accelerators can do to ease relocation costs next.
@@
Linear Probability Model: (1) (2) (3) (4) Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) Two Millions Five Millions Two Millions Five Millions Female Founder ($\alpha_{1}$) -0.113** -0.077* -0.335 -0.344*** (0.057) (0.045) (0.212) (0.117) Female×$\times$ Startup Relocated ($\alpha_{2}$) -0.062 0.015 0.265 0.410* (0.077) (0.045) (0.255) (0.229) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.175** -0.062 -0.070 -0.067 (0.085) (0.064) (0.260) (0.229) Sample Average Age in [28-40] years $N$ 337 337 337 337 $R^{2}$ 0.086 0.077 0.109 0.132 * – $p<0.1$; ** – $p<0.05$; *** – $p<0.01$
  • Notes: This table presents linear probability estimates from (3) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Only startup whose founders’ average age is between 28 and 40 are considered. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
Can accelerators level the gap? We leverage variation across accelerators to examine whether higher-quality accelerators either exhibit a smaller gender gap or help close it through the relocation channel.
@@
These results suggest that one key advantage of larger cohorts is that by engaging with a broader group of founders, relocating female entrepreneurs build a support network that fosters their startups’ growth. Further research could explore the underlying mechanisms driving this effect. For instance, female founders may not only exchange knowledge on managing a startup but also on balancing entrepreneurial and household responsibilities. Additionally, a stronger network may allow them to overcome time constraints (as the coefficients on the first row remain negative) by providing access to trusted individuals who can assist with specific business challenges. Enhanced networking opportunities may also facilitate strategic partnerships between startups or among key individuals within the ecosystem.
The remaining columns of Table 9 examine a sample of more experienced accelerators. While the gender gap persists for one-year funding milestones (columns 5 and 6), it fully disappears over the five-year horizon (columns 7 and 8). Experienced accelerators may offer structured mentorship programs tailored to female founders, equipping them with the skills necessary to successfully scale their startups.

### v2

@@ 5 Results
We estimate a two-stage model to examine how startup and accelerator characteristics influence both the matching process and subsequent performance outcomes. The first stage models the formation of matches between accelerators and startups. The second stage analyzes post-match outcomes, correcting for endogeneity that arises due to unobserved match quality.
5.1 First Stage: Matching Model
We specify the following linear utility function:
$U_{as}$ $=\beta_{1}\text{Female Founder}_{s}+\beta_{2}\text{``All Men'' Accelerator}_{a}+\beta_{3}\text{Startup Relocated}_{as}$ (5)
$\quad+\beta_{4}(\text{Female Founder}_{s}\times\text{``All Men'' Accelerator}_{a})+\bm{Z}^{as}\bm{\beta}_{Z}+\epsilon^{as},$
[…]
We estimate the parameters in equation (5) via maximum likelihood explained in Section 4.2.1. The computational details are in Appendix A.5. After obtaining parameter estimates $\hat{\bm{\beta}}$, we compute the control function term $\widehat{\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]}$, which captures the conditional expectation of the unobserved quality of a given match, given the realized matching $\mu$ and the vector of observables $\bm{X}$, which spans $\bm{X}^{as}$ for all agents in the market. [^11] Note that $\bm{X}^{as}\subset\bm{Z}^{as}$ because $\bm{X}^{as}$ includes all right-hand side variables in (5).
Appendix A.6 reports the estimation result. The model fits the data well, explaining 78% of the observed and unobserved variation in match qualities. In this section, instead, we focus on the second-stage estimates and the gender gap.
5.2 Second Stage: Outcome Equation
In the second stage, we estimate the effect of startup and accelerator characteristics on a post-match outcome $Y^{as}$, such as whether a startup secures VC funding beyond a given threshold. The estimation equation is:
$Y^{as}=$ $\alpha_{1}\text{Female Founder}_{s}+\alpha_{2}\text{``All Men'' Accelerator}_{a}+\alpha_{3}\text{Startup Relocated}_{as}$ (6)
$+\alpha_{4}(\text{Female Founder}_{s}\times\text{``All Men'' Accelerator}_{a})+\bm{Z}^{as}\bm{\alpha}_{Z}$
$+\lambda\,\widehat{\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]}+\gamma_{t}+\theta_{k}+\delta^{as},$
where $\gamma_{t}$ are year fixed effects capturing year-specific funding environments, $\theta_{k}$ are industry fixed effects based on the startup’s sector, and $\delta^{as}$ is an idiosyncratic error term. The control function term $\widehat{\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]}$ is estimated as the residual of (5) and corrects for endogeneity in the matching process, as justified by Proposition 2.
Fixed effects.
In the second stage, we include year fixed effects $\gamma_{t}$ to account for time-varying funding conditions. These are excluded from the first stage because each market is defined over a six-month cohort, which is more granular than a calendar year, and thus yearly fixed effects cannot be estimated in the first stage.
Industry fixed effects $\theta_{k}$ are included to address concerns that post-match outcomes may differ across tech sub-sectors with varying investment timelines. We do not include these fixed effects in the first stage because accelerators in our sample recruit startups across a wide range of tech industries and generally do not specialize within narrow sub-sectors. Moreover, since accelerators often hold equity stakes for several funding rounds before exit, industry-specific investment cycles are more relevant to performance outcomes than to the initial match formation. Thus, including industry fixed effects in the outcome stage helps absorb any remaining heterogeneity in investment trajectories.
@@ 5.3 Post-Accelerator Performance
The first four columns of Table 5 present selected coefficients from estimating (6) using startups’ VC fundraising performance one year after the demo day as dependent variables.
We begin with the one-year horizon because such funding is typically initiated during or shortly after the demo day, as the process from signing the initial contract to receiving the investment often takes several months. [^12] https://www.forbes.com/sites/alejandrocremades/2019/01/03/how-long-it-takes-to-raise-capital-for-a-startup/?sh=1129c4a37a41 Within this short period, the quality of startups changes minimally following graduation, providing us with sufficient control over the startup qualities that are observable to accelerators and investors but not to researchers.
@@
The results in columns (1) and (2) show that the coefficients on “Female Founder” are not statistically significant, suggesting no significant gender differences in the likelihood of receiving any VC funding or amounts below two million US\$. However, in columns (3) and (4), the coefficients on “Female Founder” become significantly negative, indicating that women-founded startups are significantly less likely to raise larger amounts of funding.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Funding Within One Year Since Graduation Funding Within Five Years Since Graduation Funded? Funding Exceeds (0/1) Funded? Funding Exceeds (0/1) (0/1) One Million Two Millions Five Millions (0/1) One Million Two Millions Five Millions Female Founder 0.053 -0.049 -0.082 -0.038 -0.029 -0.156 -0.109 -0.160 (0.130) (0.102) (0.029) (0.021) (0.123) (0.124) (0.116) (0.073) ”All Men” Accelerator 0.049 0.051 -0.023 -0.009 0.085 0.055 0.051 0.007 (0.062) (0.052) (0.034) (0.025) (0.064) (0.067) (0.061) (0.053) Female Founder $\times$ ”All Men” Accelerator -0.097 0.068 0.066 0.017 -0.097 0.168 0.068 0.103 (0.149) (0.113) (0.039) (0.024) (0.145) (0.143) (0.133) (0.090) Startup Relocated -0.068 -0.105 -0.043 -0.033 -0.046 -0.084 -0.081 -0.034 (0.052) (0.038) (0.022) (0.016) (0.057) (0.053) (0.045) (0.042) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $N$ 736 736 736 736 736 736 736 736 $R^{2}$ 0.093 0.117 0.055 0.034 0.097 0.087 0.056 0.048 Notes: This table presents linear probability estimates from (4), where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
@@
Across columns, we also examine whether “all-men” accelerators further decrease the chances of female entrepreneurs obtaining funding, as prior studies have shown that female investors do not display bias against female entrepreneurs and can help mitigate the gender gap (Raina, , 2021; Hebert, , 2023). The coefficients of “Female Founder $\times$×All Men’ Accelerator” are generally not significant for the probability of being funded (columns 1 and 5). However, it is surprisingly positive and significantive in column (3) for the probability of receiving funding of more than 2 million US\$. Taken together, this evidence does not support the idea that startups with women founders are more likely to match with accelerators with a mixed (or all-women) gender composition.
Finally, the last row of Table 8 highlights the significant costs associated with relocation. Moving to a new state reduces the probability of reaching the one-million and five-million funding milestones by approximately 3.3 to 10 percentage points. To the extent that startups securing five million in funding are of higher quality, their superiority may help them better absorb these relocation costs. Over the five-year horizon, however, the negative impact of relocation diminishes substantially.
@@ 6 What Drives the Gender Gap in Funding?
To view the role of relocating challenges, Table 6 interacts “Female Founder” and the indicator “Startup Relocated,” which is 1 if the startup relocated to a new state and zero otherwise. As in the previous table, the first four columns report startups’ performance one year after graduation, while the remaining columns extend the analysis to five years. Importantly, because of our control function approach (Proposition 2) we can consider startup relocation as exogenous as relocation was controlled for in the matching utilities (1)—as for all the other covariates in (4)—ensuring a causal interpretation of our findings.
Across columns, the estimated coefficients for “Female Founder” closely align with those in Table 5. For example, female-founded teams have a 7.6 percentage point lower probability of securing at least two million US\$ in the first year after graduation (Column 3) and a 17.9 percentage point lower probability of raising five million US\$ within five years (Column 8). However, this gap narrows for startups willing to relocate, as the coefficient in the second row is generally positive despite large standard errors due to the small sample. This result suggests that relocation mitigates the gender gap in funding.
@@
Therefore, this null result provides suggestive evidence that of the mechanism underlying the gender gap in Table 5: not all startups may be able to match with their favorite accelerator, leading to suboptimal performance and inefficiencies.
@@
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Funding Within One Year Funding Within Five Years Funded? Funding Exceeds (0/1) Funded? Funding Exceeds (0/1) (0/1) One Mil. Two Mil. Five Mil. (0/1) One Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.014 -0.078 -0.076 -0.041 -0.087 -0.195 -0.111 -0.179 (0.135) (0.101) (0.030) (0.023) (0.132) (0.129) (0.126) (0.086) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.238 0.102 -0.022 0.010 0.209 0.139 0.007 0.065 (0.143) (0.116) (0.034) (0.019) (0.143) (0.146) (0.129) (0.106) Other Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$ 0.225 0.024 -0.098 -0.031 0.122 -0.056 -0.104 -0.113 (0.149) (0.150) (0.039) (0.029) (0.144) (0.164) (0.140) (0.103) $N$ 736 736 736 736 736 736 736 736 $R^{2}$ 0.097 0.118 0.055 0.034 0.100 0.088 0.056 0.048 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
Confounding factors.
Higher-quality startups may be more likely to relocate, raising the concern that our analysis captures unobserved aspects of startup quality correlated with female-founded teams rather than the actual match quality and complementarity between startups and accelerators, despite our control function approach.
@@ 6.1 Mechanism
Costs for mother founders. Relocation may be particularly costly for mother entrepreneurs. Relocating with a newborn or a young child may be particularly difficult, regardless of the quality of the startup.
Unfortunately, we do not observe whether a founder is a mother. Moreover, controlling for parenthood alone may be insufficient, as plans to have children can also constrain mobility. While some administrative data might capture parenthood and surveys could reveal fertility intentions, these variables are unavailable in most datasets. Thus, to explore whether the gender gap is more pronounced in cases where family constraints are more likely, we restrict the sample to startups whose teams have an average age between 28 and 40. This range corresponds to typical childbearing ages among highly educated women: the median age at first birth for degree-holding women is around 28, and tends to be even higher for those in STEM fields—relevant for our sample of startup founders (Livingston, , 2015; Lappegard et al., , 2020; Schweizer and Guzzo, , 2020).
We present the results in Table 7, which focuses on the probability of observing fundings exceeding two and five million US\$ at either horizon, as the previous analyses highlight that the gender gap is especially pronounced for these funding levels, despite larger standard errors arising from the smaller sample size.
The results indicate a larger gender gap (row 1) compared to the estimates obtained from the full sample. The gender gap appears particularly large over the five-year horizon (Columns 3 and 4). For instance, female founders who do not relocate are 34 percentage points less likely to reach both the two- (non-significant) and the five-million funding milestones. This result comes as no surprise, as parenting may take significant resources from mother entrepreneurs, and the time they allocate away from their startups has a substantial economic cost five years after graduation.
The second line offers new insights into the benefits of relocation. While no clear advantage emerges over short horizons (columns 1 and 2), relocation is associated with substantial gains over the five-year horizon. Relocating female founders are 26.5 percentage points more likely to exceed the two-million funding threshold and 41 percentage points more likely to surpass five million. The latter effect is significant at the 10% level, and the point estimate exceeds the corresponding gender gap in the first row—suggesting that relocation can potentially offset the gender gap.
Notably, the fact that the largest effect appears at the five-million threshold, and only over five years, suggests that it may take time for female-founded startups in this age bracket to catch up—possibly due to support mechanisms implemented by the accelerator. In the next section, we examine how accelerators could help ease relocation costs.
@@
Linear Probability Model: (1) (2) (3) (4) Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) Two Millions Five Millions Two Millions Five Millions Female Founder ($\alpha_{1}$) -0.113 -0.077 -0.335 -0.344 (0.057) (0.045) (0.212) (0.117) Female×$\times$ Startup Relocated ($\alpha_{2}$) -0.062 0.015 0.265 0.410 (0.077) (0.045) (0.255) (0.229) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.175 -0.062 -0.070 -0.067 (0.085) (0.064) (0.260) (0.229) Sample Average Age in [28-40] years $N$ 337 337 337 337 $R^{2}$ 0.086 0.077 0.109 0.132 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Only startup whose founders’ average age is between 28 and 40 are considered. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
Can accelerators level the gap? We leverage variation across accelerators to examine the gender gap at higher-quality accelerators.
@@
These results suggest that one key advantage of larger cohorts is that by engaging with a broader group of founders, relocating female entrepreneurs build a support network that fosters their startups’ growth. Further research could explore the underlying mechanisms driving this effect. Additionally, a stronger network may allow them to overcome time constraints (as the coefficients on the first row remain negative) by providing access to trusted individuals who can assist with specific business challenges. Enhanced networking opportunities may also facilitate strategic partnerships between startups or among key individuals within the ecosystem.
The remaining columns of Table 8 examine a sample of more experienced accelerators. While the gender gap persists for one-year funding milestones (columns 5 and 6), it fully disappears over the five-year horizon (columns 7 and 8). As the investment horizons of accelerator is approximately five years, experienced accelerators may offer structured mentorship programs tailored to female founders, equipping them with the skills necessary to successfully scale their startups and catch up by this date (Zhang, , 2023).